# ReliableEval: A Recipe for Stochastic LLM Evaluation via Method of Moments

### v1

@@ Abstract
LLMs are highly sensitive to prompt phrasing, yet standard benchmarks typically report performance using a single prompt, raising concerns about the reliability of such evaluations. In this work, we argue for a stochastic method of moments evaluation over the space of meaning-preserving prompt perturbations. We introduce a formal definition of reliable evaluation that accounts for prompt sensitivity, and suggest ReliableEval – a method for estimating the number of prompt resamplings needed to obtain meaningful results. Using our framework, we stochastically evaluate five frontier LLMs and find that even top-performing models like GPT-4o and Claude-3.7-Sonnet exhibit substantial prompt sensitivity. Our approach is model-, task-, and metric-agnostic, offering a recipe for meaningful and robust LLM evaluation.
@@ 1 Introduction
We hope that our recommendations will be adopted to achieve meaningful and reliable reporting of LLM performance. [^1] Code and data available at https://github.com/SLAB-NLP/Reliable-Eval
@@ 2.1 Characterizing LLM Performance Using Distributional Analysis
$\mu_{1}(M,S_{D})$ $=\sum_{s\in S_{D}}\varepsilon_{M}(s)\cdot P(S=s)$ (2)
[…]
@@
$\mu_{2}(M,S_{D})=\frac{1}{\big{|}S_{D}\big{|}}\sum_{s\in S_{D}}\left(\varepsilon_{M}(s)-\mu_{1}\right)^{2}$ (3)
@@ 3 ReliableEval: Recipe for Stochastic Evaluation
Identify dimensions of meaning-preserving prompt perturbations that may influence model performance – such as instruction phrasing, output format, or few-shot examples. We recommend leveraging existing work aligned with the task type. For instance, for multiple-choice QA datasets, the framework by Habba et al. (2025) can be used to generate the prompt perturbation space $S_{D}$. Their approach builds on the Unitxt framework for structured data preparation, which can also be extended to generate prompt perturbations for other task types Bandel et al. (2024).
@@
Next, construct the $\delta$-level confidence interval (CI) over $\Delta(n)$, which filters $\Delta(n)$ to the range between the $\delta/2$ and $21-\delta/$ 2 percentiles. For instance, if $\delta=0.1$, the corresponding $\mathbf{CI}_{0.1}(\Delta(n))$ includes all values of $\Delta(n)$ which lie between the 5th and 95th percentiles. Then, define $n^{*}$ as the smallest $n$ for which $\epsilon$ is larger than the maximum of this confidence interval:
@@ 4.1 Experimental Setup
We run stochastic evaluation on MMLU, GPQA-Diamond, and SimpleQA. The curation of the meaning-preserving prompt perturbations space is done by leveraging unitxt Bandel et al. (2024) and Dove Habba et al. (2025).We evaluate five LLMs: Llama-3.3-70B Grattafiori et al. (2024), Deepseek-v3 Liu et al. (2024a), GPT-4o Hurst et al. (2024), Claude-3.7-Sonnet Anthropic (2025), and Grok-3 xAI (2025). As defined in Section 2, we set the following parameters to estimate a reliable evaluation $\epsilon=0.01,\>\delta=0.1,\>N=100$, with Llama-3.3-70B serving as the reference model $\hat{M}$ for estimating $n^{*}$. See additional implementation details in the Appendix.
@@ 6 Conclusion
We propose to estimate model performance over prompt variations using moment analysis and show how to compute how many samples are needed for reliable results. Evaluating frontier models across benchmarks, we find that sensitivity varies widely, underscoring the need for more robust evaluation practices.

### v2

@@ Abstract
LLMs are highly sensitive to prompt phrasing, yet standard benchmarks typically report performance using a single prompt, raising concerns about the reliability of such evaluations. In this work, we argue for a stochastic method of moments evaluation over the space of meaning-preserving prompt perturbations. We introduce a formal definition of reliable evaluation that accounts for prompt sensitivity, and suggest ReliableEval – a method for estimating the number of prompt resamplings needed to obtain meaningful results. Using our framework, we stochastically evaluate five frontier LLMs and find that even top-performing models like GPT-4o and Claude-3.7-Sonnet exhibit substantial prompt sensitivity. Our approach is model-, task-, and metric-agnostic, offering a recipe for meaningful and robust LLM evaluation. [^1] Code and data available at https://github.com/SLAB-NLP/Reliable-Eval
@@ 1 Introduction
We hope that our recommendations will be adopted to achieve meaningful and reliable reporting of LLM performance.
@@ 2.1 Characterizing LLM Performance Using Distributional Analysis
$\mu_{1}(M,S_{D})$ $=\underset{s\overset{\text{i.i.d.}}{\sim}S_{D}}{\mathbb{E}}[\varepsilon_{M}]$ (2)
$=\sum_{s\in S_{D}}\varepsilon_{M}(s)\cdot P(S=s)$
$\overset{\text{{uniform i.i.d.}}}{=}\frac{1}{\big{|}S_{D}\big{|}}\sum_{s\in S_{D}}\varepsilon_{M}(s)$
@@
$\mu_{2}(M,S_{D})$ $=\mathbb{E}_{s\overset{\text{i.i.d.}}{\sim}S_{D}}[{\varepsilon_{M}}^{2}]$ (3)
$=\mathbb{E}_{s\overset{\text{i.i.d.}}{\sim}S_{D}}[(\varepsilon_{M}(s)-\mu_{1})^{2}]$
$\overset{\text{{uniform i.i.d.}}}{=}\frac{1}{\big{|}S_{D}\big{|}}\sum_{s\in S_{D}}\left(\varepsilon_{M}(s)-\mu_{1}\right)^{2}$
@@ 3 ReliableEval: Recipe for Stochastic Evaluation
Identify dimensions of meaning-preserving prompt perturbations that may influence model performance – such as instruction phrasing, output format, or few-shot examples. We recommend leveraging existing work aligned with the task type. For instance, for multiple-choice QA datasets, the framework by Habba et al. (2025) can be used to generate the prompt perturbation space $S_{D}$. Their approach builds on the Unitxt framework for structured data preparation, which can also be extended to generate prompt perturbations for other task types Bandel et al. (2024). Notably, our proposed method is flexible and not restricted to any predefined set of meaning-preserving prompt perturbations, and other paraphrases can be used to construct the sample space $S_{D}$.
@@
Next, construct the $\delta$-level confidence interval (CI) over $\Delta(n)$, which filters $\Delta(n)$ to the range between the $\delta/2$ and $1-\delta/2$ percentiles. For instance, if $\delta=0.1$, the corresponding $\mathbf{CI}_{0.1}(\Delta(n))$ includes all values of $\Delta(n)$ which lie between the 5th and 95th percentiles. Then, define $n^{*}$ as the smallest $n$ for which $\epsilon$ is larger than the maximum of this confidence interval:
@@
We note that in some scenarios, such as when the focus is on evaluating a single model or when the variations between models is large, it may be preferable to use a reference dataset instead of a reference model. For example, if we want to evaluate model $M$ on multiple datasets, we can choose a reference dataset $D\textquoteright$, compute its empirical moments over large $N$ as a proxy for the true moments.
@@ 4.1 Experimental Setup
We run ReliableEval on MMLU Hendrycks et al. (2021), GPQA-Diamond Rein et al. (2024), and SimpleQA Wei et al. (2024), which are all widely-used English benchmarks. The curation of the meaning-preserving prompt perturbations space is done by leveraging unitxt Bandel et al. (2024) and Dove Habba et al. (2025).We evaluate five LLMs: Llama-3.3-70B Grattafiori et al. (2024), Deepseek-v3 Liu et al. (2024a), GPT-4o Hurst et al. (2024), Claude-3.7-Sonnet Anthropic (2025), and Grok-3 xAI (2025). As defined in Section 2, we set the following parameters to estimate a reliable evaluation $\epsilon=0.01,\>\delta=0.1,\>N=100$, with Llama-3.3-70B serving as the reference model $\hat{M}$ for estimating $n^{*}$. See additional implementation details in the Appendix.
@@ 6 Conclusion
We propose to estimate model performance over prompt variations using moment analysis and show how to compute how many samples are needed for reliable results.
Our proposed method is designed to accommodate any computational budget, with an inherent trade-off between budget, error margin, and confidence. The practical question becomes: “Given a specific compute budget, what is the most reliable evaluation achievable?” In our framework, the compute budget sets the maximum feasible $N$ and constrains $n^{*}$. If for a given error margin $\epsilon$ and confidence level $\delta$ the number of samples nn* exceeds a given budget, it is still possible run the evaluation with $n<n^{*}$ resamplings, accepting a larger margin of error or lower confidence as a result. Thus, even with limited compute resources, our method provides guidance on how to maximize evaluation reliability within those constraints.
Finally, by evaluating frontier models across benchmarks, we find that sensitivity varies widely, underscoring the need for more robust evaluation practices.
@@
Acknowledgments
This work was partially supported by research grant no. 7256 from the Israeli Ministry of Science and Technology. We thank Dr. Arie Cattan and Dr. Ori Shapira for the helpful discussions and advice on this project.